A 200g hockey puck is launched up a metal ramp that is inclined at a 30 degree angle. the coefficients of static and kinetic between the hockey puck and the metal ramp are Us= 0.40 and Uk=0.30. the puck's initial speed is 63m/s.

What speed does the puck have when it slides back to starting point?
Answer #1

Here ,

acceleration the puck while going up ,

a = g * sin(theta) + u * g * cos(theta)

a = 9.8 *(sin(30) + 0.5 * cos(30))

a = 7.45 m/s^2

Now , distance it moves up

d = v^2/(2 *a)

d = 63^2/(2 * 7.45)

d =270.8 m

Now, using work energy theorum ,

let the final velocity of puck is v

0.5 * m * (v^2 - 63^2) = work done by friction

0.5 * m * (v^2 - 63^2) = -2 * 270.8 * 0.30 * m * cos(30) * 9.8

solving for v

v = 50.9 m/s

the speed of puck when it come back to stating point is 50.9 m/s